# Angular momentum of vacuum bubbles in a first-order phase transition

### v1

@@ 4.2.1 Background evolution
$c_{s}^{2}$ Average sound speed squared, weighted over the dark and visible sectors
@@
$\delta$ Density contrast, weighted over the dark and visible sectors
@@
We emphasize that $c_{s,i}^{2}$ in this limit refers to the sound speed in the FV or the TV, when each is taken in isolation. This is the sound speed that has been computed in e.g. [25, 27]. However, during the phase transition, the sound speed $c_{s\rm,D}^{2}$ in the dark plasma may deviate from the value suggested by $c_{s\rm,FV}^{2}$ or $c_{s\rm,TV}^{2}$. For example, if the false vacuum fraction decreases sharply, the denominator in Eq. (4.45) can be large so that $c_{s\rm,D}^{2}$ can be close to zero. A similar situation occurs in the physical system considered in Ref. [21], where the system undergoes prolonged phase coexistence at $T\simeq T_{\rm c}$, keeping the pressure constant, which leads to a zero sound speed in the long-wavelength limit. On the other hand, Ref. [27] mentioned in a footnote that the sound speed does not vanish at the critical point, which is contrary to the result in Ref. [21]. This is not very surprising, since the former considered the sound speeds in each phase taken in isolation, i.e. $c_{s,\rm FV}^{2}$ and $c_{s\rm,TV}^{2}$. For the sake of clarity, in Table 2 we explicitly define those quantities, such as the equation of state and sound speed squared, that refer to: the different phases in the dark sector; the weighted average over phases in the dark sector; the weighted average over fluids in the cosmological system.
Like the background evolution history, the evolution of the FV equation of state and the sound speed qualitatively differ between the two benchmark cases in Fig. 3. Based off Eq. (2.21), the sound speed for BP-2, with a higher percolation temperature, plateaus to a larger $c_{s\rm,D}^{2}\simeq c_{s\rm,TV}^{2}$ than BP-7, shown as solid curves in the bottom row of Fig. 3. On the other hand, we both see dip features in the $c_{s\rm,D}^{2}$ evolution. We can estimate the duration $\eta_{f}-\eta_{i}$ when the dark plasma sound speed deviates significantly from either $c_{s\rm,FV}^{2}$ and $c_{s\rm,TV}^{2}$ as follows. From Eq. (4.41) the rate of change in FV fraction is
[…]
Since $\dot{T}=-HT$, the temperature decreases monotonically with time, and we take $T_{i}=T(\eta_{i})$ to be the temperature when $\dot{F}=-HF$, i.e. the relative decrease in the FV fraction within a Hubble time is order unity. Then
$8\pi\frac{v_{\rm w}^{3}(T_{i})}{(\beta/H)^{3}(T_{i})}\frac{\Gamma(T_{i})}{H^{4}(T_{i})}=1\Rightarrow 8\pi\frac{v_{\rm w}^{3}(T_{i})}{(\beta/H)^{3}(T_{i})}\frac{\Gamma(T_{i})}{H_{\rm c}^{4}}\left(\frac{T_{\rm c}}{T_{i}}\right)^{8}\simeq 1\,.$ (4.48)
In the last equation, we approximated the Hubble parameter as
@@
The first step assumes that the SM dominates the Hubble parameter. In the last step we used the fact that the dark-to-SM temperature ratio is constant since both $T$ and $T_{\rm SM}$ scale as the inverse of the scale factor. On the other hand, $T_{f}=T(\eta_{f})$ is taken to be the moment when $F\ll 1$; we then set $-\ln F(\eta_{f})=10$, and thus
$8\pi\frac{v_{\rm w}^{3}(T_{f})}{(\beta/H)^{4}(T_{f})}\frac{\Gamma(T_{i})}{H^{4}(T_{f})}=10\Rightarrow 8\pi\frac{v_{\rm w}^{3}(T_{f})}{(\beta/H)^{4}(T_{f})}\frac{\Gamma(T_{f})}{H_{\rm c}^{4}}\left(\frac{T_{\rm c}}{T_{f}}\right)^{8}\simeq 10\,.$ (4.50)
Then the time duration in which the dark sound speed exhibits a dip is
In the bottom panels of Fig. 3 we indicated the estimates for $\eta_{i}$ and $\eta_{f}$ as vertical dashed and dotted lines, respectively, and overlaid them on the $c_{s\rm,D}^{2}$ curves. For these two BPs we can clearly see that our definitions of $\eta_{i}$ and $\eta_{f}$ provide us with a robust method to determine the width of the dip in $c_{s\rm,D}^{2}$. As for the equation of state, note that $w_{\rm D}$ changes with conformal time via
@@ 4.2.2 Parameter selection and scans over physical FOPT scenarios
We have shown in Section 2.1 that the physical conditions restrict the allowed regions in the $(L_{\rm c}/T_{\rm c}^{4}-x_{C})$ plane, for given $A$ and $\lambda$. We observed that for a given $A$, a larger $\lambda$ allows part of the physical region to reach values of $T_{0}/T_{\rm c}$ close to 1. Since the destabilization temperature $T_{0}$ determines the lower limit of the percolation temperature $T_{*}$, there are situations where $T_{*}$ is guaranteed to be extremely close to $T_{\rm c}$. We would like to avoid such scenarios, which correspond to small potential barriers.
Refer to caption Refer to caption
(a) (b)
Figure 5: Panel (a) shows the allowed range of $L_{\rm c}/T_{\rm c}^{4}$ values versus $\lambda$, for each $A$ value. Panel (b) shows the allowed range of $T_{0}/T_{\rm c}$, for each $A$ value. In both cases, $x_{C}=0$; the vertical dashed lines indicate $\lambda_{0}(A)$.
In Fig. 5 we plot, for a given value of $A$, the maximum range of $L_{\rm c}/T_{\rm c}^{4}$ and $T_{0}/T_{\rm c}$ values as functions of $\lambda$, which is achieved by setting $x_{C}=0$. For sufficiently large $\lambda$, the maximum $T_{0}/T_{\rm c}$ approaches unity; correspondingly, the maximum value of $L_{\rm c}/T_{\rm c}^{4}$ is set by $g_{\rho}\pi^{2}/30\simeq 1.48$. Furthermore, setting the maximum $T_{0}/T_{\rm c}=0.95,0.99,0.995,0.999$ corresponds, respectively, to $\lambda/\lambda_{0}\simeq 1.42,2.34,2.94,5$, for $A=10^{-1},10^{-2},10^{-3},10^{-4}$. The actual values of $\lambda$ for each $A$ are shown in Table 3, which provides us with a guide to select the benchmark points in Table 1.
$A$ $\lambda_{0}$ $\lambda_{0.95}$ $\lambda_{0.99}$ $\lambda_{0.995}$ $\lambda_{0.999}$
$10^{-1}$ 0.129 0.183 0.303 0.380 0.647
$10^{-2}$ 0.00600 0.00850 0.0140 0.0176 0.0300
$10^{-3}$ 0.000279 0.000395 0.000652 0.000818 0.00139
$10^{-4}$ $1.29\times 10^{-5}$ $1.83\times 10^{-5}$ $3.03\times 10^{-5}$ $3.80\times 10^{-5}$ $6.47\times 10^{-5}$
Table 3: For each $A$, we determine the $\lambda$ values where $T_{0}/T_{\rm c}=0.95,0.99,0.995,0.999$ for $x_{C}=0$. These are respectively denoted by $\lambda_{0.95}$, …, $\lambda_{0.999}$. Note that for each $A$, the ratios $\lambda_{0.95}/\lambda_{0},...,\lambda_{0.999}/\lambda_{0}$ are respectively given by 1.416, 2.340, 2.936, and 5.004.
@@
Figure 6: Contour plots of $T_{*}/T_{\rm c}$ for $A=10^{-1},10^{-2},10^{-3},10^{-4}$, for fixed $\lambda/\lambda_{0}=1.42$, and $T_{\rm c}=10\,\mathrm{MeV}$, $r_{T\rm,c}=0.4$.
Restricting our attention to physical FOPT scenarios, we would like to determine how the fundamental parameters that characterize the FOPT determine physically relevant quantities, such as the percolation temperature and the strength of the phase transition. This requires us to specify $T_{\rm c}$ and $r_{T\rm,c}$, and then solve the system of ODEs that describe the background evolution. Firstly, in Fig. 6 we show scans of $T_{*}/T_{\rm c}$. Here we take $A=10^{-1},10^{-2},10^{-3},10^{-4}$. For each $A$, we choose $\lambda=0.303,1.4\times 10^{-2},6.52\times 10^{-4},3.03\times 10^{-5}$, respectively, i.e. we fix $\lambda/\lambda_{0}=1.42$. We commit to $T_{\rm c}=10\,\mathrm{MeV}$ and $r_{T\rm,c}=0.4$, unless we specify otherwise. For a fixed $x_{C}$, we find that the percolation temperature $T_{*}$ tends to become closer to the critical temperature, along the direction of rescaled latent heat. This is expected behavior, knowing that $T_{0}$ also approaches $T_{\rm c}$ along this direction (see Fig. 1). We have checked that the general trend on $T_{*}$ still holds for different $T_{\rm c}$ and $r_{T\rm,c}$; also, the contours for fixed $A$ and $\lambda/\lambda_{0}$ change only marginally. Thus, we find that imposing the condition that $T_{*}$ is at most $0.95T_{\rm c}$ restricts us from going close to the edge of the physical region towards the maximum rescaled latent heat.
@@
Figure 7: Contour plots of $\alpha_{*}$ for $A=10^{-1},10^{-2},10^{-3},10^{-4}$, for fixed $\lambda/\lambda_{0}=1.42$, and $T_{\rm c}=10\,\mathrm{MeV}$, $r_{T\rm,c}=0.4$.
In Fig. 7 we show scans of the FOPT strength, defined in Eq. (4.39). To understand the scans, we estimate $\alpha_{*}$ by noting that we are working in the regime where $\rho_{\rm SM}\gg\rho_{\rm D}$. This is required by the limit on
$\Delta N_{\rm eff}=\frac{4}{7}\left(\frac{11}{4}\right)^{4/3}g_{\rho}r_{T\rm,c}^{4}\simeq 0.25\left(\frac{g_{\rho}}{4.5}\right)\left(\frac{r_{T\rm,c}}{0.4}\right)^{4}\leq 0.5\,,$ (4.54)
from a combination of Planck and other data [44], which translates to $r_{T\rm,c}<0.474$. Then,
$\rho_{\rm tot}\approx\rho_{\rm SM}\approx\frac{\pi^{2}}{30}\frac{g_{\rho,{\rm SM}}}{r_{T\rm,c}^{4}}T_{*}^{4}\,,$ (4.55)
which gives
$\alpha_{*}\approx\frac{1}{(T_{*}/T_{\rm c})^{4}}\left(1-\frac{3}{4}\frac{T_{*}}{T_{\rm c}}\right)\frac{L_{\rm c}}{T_{\rm c}^{4}}r_{T\rm,c}^{4}\,.$ (4.56)
The trend in percolation temperature in Fig. 6, combined with the estimate for $\alpha_{*}$ in Eq. (4.56), are consistent with the scans of $\alpha_{*}$ in Fig. 7.
4.2.3 Perturbations during a FOPT
Once we know the background evolution during the FOPT, we can determine the evolution of each Fourier mode $k/\mathcal{H}_{\rm c}$ by solving Eqs. (4.3) and (4.4). In the following discussion, we assume that the single-fluid picture of the dark plasma holds for the perturbation modes of interest. The expressions for the average pressure, entropy density and energy density are provided in Eq. (4.44), and those for the equation of state and sound speed in Eq. (4.45).
@@ 4.2.3 Perturbations during a FOPT
$\Phi_{k}\simeq\frac{3}{2}\left(\frac{k}{\mathcal{H}}\right)^{-2}\delta_{\text{SM},k}\,.$ (4.66)
Figure 8: Comparison between exact $\delta$ and $\hat{\psi}$ evolution versus numerically calculated SM perturbations.
We first examine the behavior of the SM perturbations. In Fig. 8 we compare the numerical solutions for $\delta_{\rm SM}$ and $\hat{\psi}_{\rm SM}$ from solving the UHG perturbation equations during the FOPT, and the result that one would obtain from the analytic expressions in Eq. (4.13), that are valid during the period when the dominant component of the Universe is a pure radiation fluid. Choosing BP-2, we plot the difference between the numerical result and the analytic result, normalized with respect to the RMS value of the analytic result within the period between the critical point and the percolation time, for the modes $k/\mathcal{H}_{\rm c}=100$ and $k/\mathcal{H}_{\rm c}=1000$. The difference is negligible, mainly because the equation of state and sound speed in the SM plasma are unaffected by the FOPT in the dark sector. To a good approximation, the analytic expressions for the perturbations in Eq. (4.13) can be extrapolated even during the period of FOPT for this benchmark case.
@@ 4.2.4 Superhorizon and subhorizon limits
where we have dropped terms containing positive powers of $z$ and higher. We can compare these with [21], which found that the evolution of the perturbations in the superhorizon limit is independent of the sound speed. This applies in our case if the dominant fluid component, which mainly dictates the Hubble expansion, is the dark fluid. One can show that the continuity equation can be rewritten as
@@
A similar expression was found in [21], apart from the extra terms appearing in the $\delta_{\rm D}$ term. In the WKB regime, we assume that the oscillation period of the Fourier mode under consideration is much shorter than the scale of the time variation of the sound speed and of $c_{s\rm,D}^{2}-w_{\rm D}$, i.e.
$k^{2}c_{s\rm,D}^{2}\gg\mathcal{H}~{}\frac{1}{c_{s\rm,D}^{2}}\left(c_{s\rm,D}^{2}\right)^{\prime},\quad k^{2}c_{s\rm,D}^{2}\gg\mathcal{H}^{2}\left(c_{s\rm,D}^{2}-w_{\rm D}\right)\,.$ (4.77)
Following [21], we introduce
@@
where we essentially took the zeroth order coefficient to be dominated by the time dependent frequency $\omega(\eta)\equiv kc_{s\rm,D}(\eta)$. The WKB solution then becomes
@@
Armed with the WKB approximation to $u_{\rm D}$, we can use Eq. (4.78) to obtain $\delta_{\rm D}$. For a given Fourier mode, the coefficients $a_{k}$ and $b_{k}$ are chosen to match the initial conditions, which are expressed in terms of $\delta_{\rm D}$ and $\delta_{\rm D}^{\prime}$, where
$\delta_{\rm D}(\eta_{\rm c})=\delta_{\rm SM}(z_{\rm c})\,,\quad\delta_{\rm D}^{\prime}(\eta_{\rm c})=\hat{\psi}_{\rm SM}(z_{\rm c})~{}\sqrt{3}\mathcal{H}_{\rm c}z_{\rm c}\,.$ (4.81)
Meanwhile, we have
$\hat{\psi}_{\rm D}\approx\frac{\delta_{\rm D}^{\prime}}{k}\,,$ (4.82)
which we can use to obtain $\hat{\psi}_{\rm D}$.

### v2

@@ 4.2.1 Background evolution
$c_{s}^{2}$ Sound speed squared, averaged over the dark and visible sectors
@@
$\delta$ Density contrast, averaged over the dark and visible sectors
@@
Like the background evolution history, the evolution of the dark fluid equation of state and the sound speed qualitatively differ between the two benchmark cases in Fig. 3. For example, we observe that the period when $c_{s\rm,D}^{2}$ deviates significantly from $1/3$ is shorter for BP-4 than for BP-2. We can estimate the duration $\eta_{f}-\eta_{i}$ when the dark plasma sound speed deviates significantly from either $c_{s\rm,FV}^{2}$ and $c_{s\rm,TV}^{2}$ as follows. We take $\eta_{i}$ and $\eta_{f}$ to be the instants in time when the fractional rate of change in the FV fraction $\dot{F}/(HF)$, are, respectively, 1 and 10. These can be calculated numerically. However, an estimate can be obtained from Eq. (4.42) assuming that $v_{\rm w}$ and $\beta$ are slowly varying functions of $t$:
$\frac{-\dot{F}}{HF}$ $\approx 8\pi\frac{v_{\rm w}^{3}}{\beta^{3}(t)}\frac{\Gamma(t)}{H(t)}\,.$ (4.48)
The duration for which $c_{s\rm,D}^{2}$ exhibits a dip is
To estimate $T_{i}$ and $T_{f}$, we note that $v_{\rm w}$, $\beta$, $\Gamma$, and $H$ are functions of the dark sector temperature, and the Hubble parameter can be approximated as
@@
Then
$8\pi\frac{v_{\rm w}^{3}(T_{i})}{(\beta/H)^{3}(T_{i})}\frac{\Gamma(T_{i})}{H_{\rm c}^{4}}\left(\frac{T_{\rm c}}{T_{i}}\right)^{8}\simeq 1,\quad 8\pi\frac{v_{\rm w}^{3}(T_{f})}{(\beta/H)^{3}(T_{f})}\frac{\Gamma(T_{f})}{H_{\rm c}^{4}}\left(\frac{T_{\rm c}}{T_{f}}\right)^{8}\simeq 10\,.$ (4.51)
In the bottom panels of Fig. 3, we indicate the estimates for $\eta_{i}$ and $\eta_{f}$ as vertical dashed and dotted lines, respectively, and overlay them on the $c_{s\rm,D}^{2}$ curves. For these two BPs, we can clearly see that our definitions of $\eta_{i}$ and $\eta_{f}$ provide us with a robust method to determine the width of the dip in $c_{s\rm,D}^{2}$. As for the equation of state, note that $w_{\rm D}$ changes with conformal time via
@@ 4.2.2 Parameter selection and scans over physical FOPT scenarios
We have shown in Section 2.1 that the physical conditions restrict the allowed regions in the $(L_{\rm c}/T_{\rm c}^{4},x_{C})$ plane, for given $A$ and $\lambda$. We observed that for a given $A$, a larger $\lambda$ allows part of the physical region to reach values of $T_{0}/T_{\rm c}$ close to 1. Since the destabilization temperature $T_{0}$ determines the lower limit of the percolation temperature $T_{*}$, there are situations where $T_{*}$ is guaranteed to be extremely close to $T_{\rm c}$. We avoid such scenarios, which correspond to small potential barriers. The maximum possible $T_{0}/T_{\rm c}$ allowed by the physical conditions can be obtained by setting $\rho_{\rm TV}(T_{0})=0$ at $x_{C}=0$:
$\frac{\lambda}{\lambda_{0}(A)}=\left(\frac{9}{128}\right)^{1/3}\left(\frac{9T_{0}^{2}/T_{\rm c}^{2}-1}{1-T_{0}^{2}/T_{\rm c}^{2}}\right)^{1/3}.$ (4.54)
For any given value of $A$, if we set the maximum $T_{0}/T_{\rm c}=0.9,0.95,0.99,0.995,0.999$, they correspond, respectively, to $\lambda/\lambda_{0}\simeq 1.33,1.72,3.02,3.82,6.55$.
@@
Figure 5: Contours constant of $T_{*}/T_{\rm c}$ (red) for $A=10^{-3},10^{-4}$, for $\lambda/\lambda_{0}\simeq 1.33$ and $\lambda/\lambda_{0}\simeq 1.72$, and $T_{\rm c}=10\,\mathrm{MeV}$, $r_{T\rm,c}=0.4$. The blue contour marks the boundary of the physical region.
Restricting our attention to physical FOPT scenarios, we would like to determine how the fundamental parameters that characterize the FOPT determine physically relevant quantities, such as the percolation temperature and the strength of the phase transition. This requires us to specify $T_{\rm c}$ and $r_{T\rm,c}$, and then solve the system of ODEs that describe the background evolution. Firstly, in Fig. 5 we show scans of $T_{*}/T_{\rm c}$. Here we take $A=10^{-3},10^{-4}$. For each $A$, we choose $\lambda=3.6917\times 10^{-4},1.7136\times 10^{-5}$, respectively, for fixed $\lambda/\lambda_{0}\simeq 1.33$ (and $\lambda=4.806\times 10^{-4},2.231\times 10^{-5}$, respectively, for fixed $\lambda/\lambda_{0}\simeq 1.72$). We commit to $T_{\rm c}=10\,\mathrm{MeV}$ and $r_{T\rm,c}=0.4$, unless we specify otherwise. For a fixed $x_{C}$, we find that the percolation temperature $T_{*}$ tends to become closer to the critical temperature, along the direction of rescaled latent heat. This is expected behavior, knowing that $T_{0}$ also approaches $T_{\rm c}$ along this direction (see Fig. 1). We have checked that the general trend on $T_{*}$ still holds for different $T_{\rm c}$ and $r_{T\rm,c}$; also, the contours for fixed $A$ and $\lambda/\lambda_{0}$ change only marginally. Thus, we find that imposing the condition that $T_{*}$ is at most $0.95T_{\rm c}$ restricts us from going close to the edge of the physical region towards the maximum rescaled latent heat.
@@
Figure 6: Contours of constant of $\log_{10}\alpha_{*}$ (purple) in steps of 0.5 for $A=10^{-3},10^{-4}$, for fixed $\lambda/\lambda_{0}\simeq 1.33$ and $\lambda/\lambda_{0}\simeq 1.72$, and $T_{\rm c}=10\,\mathrm{MeV}$, $r_{T\rm,c}=0.4$. The blue contour marks the boundary of the physical region.
In Fig. 6 we show scans of the FOPT strength, defined in Eq. (4.39). To understand the scans, we estimate $\alpha_{*}$ by noting that we are working in the regime where $\rho_{\rm SM}\gg\rho_{\rm D}$. Then
$\rho_{\rm tot,*}\approx\rho_{\rm SM}(T_{\rm SM,*})\approx\frac{\pi^{2}}{30}\frac{g_{\rho,{\rm SM}}(T_{\rm SM,*})}{r_{T\rm,*}^{4}}T_{*}^{4}\,$, (4.55)
gives
$\alpha_{*}\approx\frac{30}{\pi^{2}g_{\rho,\rm SM}(T_{\rm SM,*})}\frac{r_{T\rm,*}^{4}}{(T_{*}/T_{\rm c})^{4}}\frac{\Delta\rho(T_{*})}{T_{\rm c}^{4}}\,$, (4.56)
where we made the approximation that the SM sector contributes dominantly to the total radiation energy density. The trend in percolation temperature in Fig. 5, combined with the estimate for $\alpha_{*}$ in Eq. (4.56), are consistent with the scans of $\alpha_{*}$ in Fig. 6. If we assume instantaneous reheating immediately after the onset of percolation, then the reheating temperature is
$\frac{\pi^{2}}{30}g_{\rho}T_{*}^{4}-(1-F_{*})\Delta\rho(T_{*})=\frac{\pi^{2}}{30}g_{\rho}T_{\rm rh}^{4}-\Delta\rho(T_{*})\Rightarrow T_{\rm rh}=T_{*}\left[1+\frac{30}{\pi^{2}g_{\rho}}F_{*}\frac{\Delta\rho(T_{*})}{T_{*}^{4}}\right]^{1/4}\,.$ (4.57)
Then applying separate entropy conservation in the SM sector and the dark sector, we find the effective number of extra massless neutrinos at the end of Big Bang Nucleosynthesis (BBN) to be
$\Delta N_{\rm eff}\equiv N_{\rm eff}-3.044=\frac{4}{7}\left(\frac{11}{4}\right)^{4/3}g_{\rho}r_{T\rm,c}^{4}\left[\frac{g_{s,\rm SM}(T_{\rm SM,BBN})}{g_{s,\rm SM}(T_{\rm SM,*})}\right]^{4/3}\left[1+\frac{30}{\pi^{2}g_{\rho}}F_{*}\frac{\Delta\rho(T_{*})}{T_{*}^{4}}\right]\leq 0.5\,$, (4.58)
where $g_{s\rm,SM}$ is the number of relativistic degrees of freedom associated with the entropy density in the SM sector, and the upper bound on $\Delta N_{\rm eff}$ arises from a combination of Planck and other data [Planck:2018vyg]. Taking BBN to end at $T_{\rm SM,BBN}=0.01\,\mathrm{MeV}$, we set $g_{s\rm,SM}\left(T_{\rm SM,BBN}\right)\simeq 3.93872$. The upper limit of the term in square brackets is $1+F_{*}$, and follows from the physical condition that $\rho_{\rm TV}(T)=g_{\rho}T^{4}~\pi^{2}/30-\Delta\rho(T)\geq 0$, which implies that $\Delta\rho(T_{*})/\left(\pi^{2}/30~g_{\rho}T_{*}^{4}\right)\leq 1$. This leads to a conservative upper limit of
$r_{T\rm,c}<0.598\left[\frac{g_{s,\rm SM}\left(T_{\rm SM,*}\right)}{10}\right]^{1/3}\left(\frac{4.5}{g_{\rho}}\right)^{1/4}\,.$ (4.59)
4.2.3 Perturbations during an FOPT
Once we know the background evolution during the FOPT, we can determine the evolution of each Fourier mode $k/\mathcal{H}_{\rm c}$ by solving Eqs. (4.3) and (4.4). In the following discussion, we assume that the single-fluid picture of the dark plasma holds for the perturbation modes of interest. The expressions for the average pressure, entropy density, and energy density, in the dark sector, are provided in Eq. (4.44); the equation of state and sound speed are given in Eq. (4.45).
@@ 4.2.3 Perturbations during an FOPT
$\Phi_{k}\simeq\frac{3}{2}\left(\frac{k}{\mathcal{H}}\right)^{-2}\delta_{{\rm SM,}k}\,.$ (4.69)
Figure 7: Comparison between exact $\delta$ and $\hat{\psi}$ evolution versus numerically calculated SM perturbations for BP-4 and $k/\mathcal{H}_{\rm c}=100,1000$.
We first examine the behavior of the SM perturbations. In Fig. 7, we compare the numerical solutions for $\delta_{\rm SM}$ and $\hat{\psi}_{\rm SM}$ from solving the UHG perturbation equations during the FOPT, and the result that one would obtain from the analytic expressions in Eq. (4.13), that are valid during the period when the dominant component of the Universe is a pure radiation fluid. Choosing BP-4, we plot the difference between the numerical result and the analytic result, normalized with respect to the RMS value of the analytic result within the period between the critical point and the percolation time, for the modes $k/\mathcal{H}_{\rm c}=100$ and $k/\mathcal{H}_{\rm c}=1000$. The difference is negligible, mainly because the equation of state and sound speed in the SM plasma are unaffected by the FOPT in the dark sector. To a good approximation, the analytic expressions for the perturbations in Eq. (4.13) can be extrapolated even during the period of FOPT for this benchmark case.
@@ 4.2.4 Superhorizon and subhorizon limits
where we have dropped terms containing positive powers of $z$ and higher. We can compare these with [Schmid:1998mx], which found that the evolution of the perturbations in the superhorizon limit is independent of the sound speed. This applies in our case if the dominant fluid component, which mainly dictates the Hubble expansion, is the dark fluid. One can show that the continuity equation can be rewritten as
@@
A similar expression was found in Ref. [Schmid:1998mx], apart from the extra terms appearing in the $\delta_{\rm D}$ term. We then consider the regime where the perturbation is deep in the subhorizon limit, so that the scale of the perturbation is much shorter than the scale of the time variation of the sound speed and of $c_{s\rm,D}^{2}-w_{\rm D}$, i.e.,
$k^{2}c_{s\rm,D}^{2}\gg\mathcal{H}~\frac{1}{c_{s\rm,D}^{2}}\left|\left(c_{s\rm,D}^{2}\right)^{\prime}\right|,\quad k^{2}c_{s\rm,D}^{2}\gg\mathcal{H}^{2}\left(c_{s\rm,D}^{2}-w_{\rm D}\right)\,.$ (4.80)
Following [Schmid:1998mx], we introduce
@@
where we essentially took the zeroth order coefficient to be dominated by the time dependent frequency $\omega(\eta)\equiv kc_{s\rm,D}(\eta)$. Then we approximate $c_{s\rm,D}^{2}$ by a piecewise linear function with
@@
$\delta=\delta_{\rm c}\cos\left(\frac{1}{\sqrt{3}}\frac{k}{\mathcal{H}_{\rm c}}\tilde{\eta}\right)+\sqrt{3}\hat{\psi}_{\rm c}\sin\left(\frac{1}{\sqrt{3}}\frac{k}{\mathcal{H}_{\rm c}}\tilde{\eta}\right),\quad\hat{\psi}=\frac{\delta^{\prime}}{k}\,$, (4.84)
where $\delta_{\rm c}$ and $\hat{\psi}_{\rm c}$ are the perturbations evaluated at the critical point, assuming the analytic expression for perturbations evolving in a pure radiation background. On the other hand, in the interval $\tilde{\eta}_{i}<\tilde{\eta}\leq\tilde{\eta}_{*}$, the ODE admits a general solution in terms of Airy functions, so that we have
$\left(\begin{matrix}\delta\\ (k/\mathcal{H}_{\rm c})\hat{\psi}\end{matrix}\right)=M(k/\mathcal{H}_{\rm c},u)\left(\begin{matrix}c_{1}\\ c_{2}\end{matrix}\right),\quad u\equiv\frac{-d_{1}-d_{2}\tilde{\eta}}{(-d_{2})^{2/3}}\,$, (4.85)
where
$d_{1}=\frac{1}{3}\left(\frac{k}{\mathcal{H}_{\rm c}}\right)^{2}\left[1-3\tilde{\eta}_{i}\left(\frac{c_{s\rm,D*}^{2}-1/3}{\tilde{\eta}_{*}-\tilde{\eta}_{i}}\right)\right]\,,\quad d_{2}=\left(\frac{k}{\mathcal{H}_{\rm c}}\right)^{2}\left(\frac{c_{s\rm,D*}^{2}-1/3}{\tilde{\eta}_{*}-\tilde{\eta}_{i}}\right)\,$, (4.86)
and
$M(k/\mathcal{H}_{\rm c},u)\equiv\left(\begin{matrix}Ai(u)&Bi(u)\\ \left(-d_{2}\right)^{1/3}Ai^{\prime}(u)&\left(-d_{2}\right)^{1/3}Bi^{\prime}(u)\end{matrix}\right)\,.$ (4.87)
The coefficients $c_{1}$ and $c_{2}$ are obtained by matching the solutions at $\tilde{\eta}=\tilde{\eta}_{i}$, which results in
$\left(\begin{matrix}c_{1}\\ c_{2}\end{matrix}\right)$ $=M^{-1}(k/\mathcal{H}_{\rm c},u_{i})\left(\begin{matrix}\delta(\tilde{\eta}_{i})\\ (k/\mathcal{H}_{\rm c})\hat{\psi}(\tilde{\eta}_{i})\\ \end{matrix}\right),$ (4.88)
$\left(\begin{matrix}\delta(\tilde{\eta}_{i})\\ (k/\mathcal{H}_{\rm c})\hat{\psi}(\tilde{\eta}_{i})\\ \end{matrix}\right)$ $=T(k)\left(\begin{matrix}\delta_{\rm c}\\ \sqrt{3}\hat{\psi}_{\rm c}\end{matrix}\right),\quad T(k)\equiv\left(\begin{matrix}\cos\left(\frac{1}{\sqrt{3}}\frac{k}{\mathcal{H}_{\rm c}}\right)&\sin\left(\frac{1}{\sqrt{3}}\frac{k}{\mathcal{H}_{\rm c}}\right)\\ -\frac{k}{\sqrt{3}\mathcal{H}_{\rm c}}\sin\left(\frac{1}{\sqrt{3}}\frac{k}{\mathcal{H}_{\rm c}}\right)&\frac{k}{\sqrt{3}\mathcal{H}_{\rm c}}\cos\left(\frac{1}{\sqrt{3}}\frac{k}{\mathcal{H}_{\rm c}}\right)\end{matrix}\right)\,.$ (4.89)